the question is .. the book says write the equation in slope intercept form and standard form .. and the equation is y - 50 = 8 (x - 4 ) not for standard form i got 8x - y = - 18 .. but for slope intercept i dont understand how you do it .. i looked in to book for te answer and it was y = 8x + 18 .. i keep foing over it but i dont understand how they got that solution
Found 2 solutions by smartypants, stanbon:
Answer by smartypants(7)

8x-y=-18
-8x -8x

-y=-18-8x
__ ______
-1 -1
y=8x+18

Answer by stanbon(75887)

y-50 =8(x-4)
Slope intercept form is y=mx+b
Solve you equation for "y" as follows:
y = 8x-32+50
y= 8x+18
Cheers,
Stan H.